OpenStudy (anonymous):

An air traffic controller spots two planes at the same altitude converging on a point as they fly at right angles to each other. One plane is 225 miles from the point moving at 450 miles per hour. The other plane is 300 miles from the point moving at 600 miles per hour.

OpenStudy (anonymous):

(a) At what rate is the distance between the planes decreasing?

OpenStudy (anonymous):

(b) How much time does the air traffic controller have to get one of the planes on a different flight path?

hero (hero):

Use a^2 + b^2 = d^2 and take the derivative with respect to time. Then find dd/dt

hero (hero):

Did you figure it out yet?

OpenStudy (anonymous):

d = sqrt ( 225^2 + 300^2) = 375 mi

OpenStudy (anonymous):

Could you take derivative of a^2 + b^2 = d^2 ?